OpenStudy (anonymous):

Seven movies (A, B, C, D, E, F, and G) are being scheduled for a showing. The order of showing is determined by random selection. Find the probability that film C will be shown first, film A next-to-last, and film E last.

OpenStudy (anonymous):

whats the formula? and which numbers go where?

OpenStudy (anonymous):

there are a couple of ways to do this

OpenStudy (anonymous):

number of ways to arrange all these 7 letters is \[7!\]

OpenStudy (anonymous):

you want to know how are favorable, that is, how many look like C, _ , _ , _ , _ , A, E

OpenStudy (anonymous):

you have 4 open slots to fill with 4 letters, and the number of ways you can do that is \(4!\)

OpenStudy (anonymous):

1/7

OpenStudy (anonymous):

so one method is to compute (if i am not mistaken) \(\frac{4!}{7!}\) now we can try it another way

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

probability that C is first is \[\frac{1}{7}\] then the probability that the second is not A or E is \[\frac{4}{6}\] the probability the next is not A or E is \[\frac{3}{5}\] then \[\frac{2}{4}\] then \[\frac{1}{3}\] finally the probability the next is A is \[\frac{1}{2}\] and these all must occur, so you can multiply them together and see if we get the same answer

OpenStudy (anonymous):

we did. i just needed to get down to 1/210

OpenStudy (anonymous):

yes, it is the same answer either method

jimthompson5910 (jim_thompson5910):

another way to do it there is only one way to pick slots 1, 6 and 7 to be C, A, and E respectively this is out of 7 P 3 = (7!)/((7-3)!) = 210 ways total to pick 3 movies (for slots 1, 6 and 7) So the probability is 1/210

jimthompson5910 (jim_thompson5910):

this trick works because you don't need to worry about the ordering of the 4 slots (2 through 6)

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

and much snappier i would say it is good to see there is more than one way to skin a cat

OpenStudy (skullpatrol):

Three lessons in one.

jimthompson5910 (jim_thompson5910):

oops meant to say the 4 slots 2 through 5, but you get the idea

OpenStudy (anonymous):

yeah i got the idea didn't think of it this way